# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

### v1

@@ 4.2 Main Results
Both CLAP and baseline methods are pre-trained on the whole training set of NuScenes [2] without labels. To compare the effectiveness of CLAP and UniPAD [53], we randomly sample 5% of the training labels and fine-tune BEVFusion [27] initialized by the pre-trained weights. Furthermore, as the size of pre-training dataset is limited, we randomly sample 2.5%, 1% and 0.5% of the training labels and fine-tune BEVFusion with CLAP pre-trained weights to see whether potential scaling property exists in pre-training with CLAP.
Comparison to Previous SOTA Method. As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $300\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [53]. For NDS metric, UniPAD [53] only achieves comparable performance while CLAP surpasses the train-from-scratch model by 1.76% . When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
Potential Scaling Property. As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning, which increases the ratio between pre-training data and fine-tuning data. The results are shown in Table 2 together with the results of fine-tuning on 5% of training set in NuScenes. It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data. These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
@@
Table 2: Results on potential scaling property of CLAP . “Init.” means the method for initialization. “Fine-tune Data” stands for the sample number of training set in NuScenes [2] used for fine-tuning/training-from-scratch. We gradually increase the number of training iterations for training-from-scratch model (BEVFusion [27]) with different numbers of training data until we observe convergence, leading to the results of “Rand.”. Then we initialize BEVFusion with weights pre-trained by CLAP and then fine-tune the model with the same numbers of iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance. All the results here are in %.
@@ 4.3 Visualizations
In this section, we provide visualization to show the effectiveness of the proposed Curvature Sampling. In order to evaluate whether the estimated curvature is able to help sampling more informative part of the environment, we use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3. The visualization results are shown in Figure 4. We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation. Furthermore, for better understanding about the curvature estimation, we zoom in for some correctly estimated regions. It can be found that though some noise exists, CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
@@ 4.4 Ablation Study
We conduct ablation study to evaluate the effectiveness of different components. As shown in Table 3, it can be found that using uniform sampling (second line) to make joint pre-training feasible does not bring improvement over separate pre-training (first line) [53]. Then, Curvature Sampling (third line) improves the performance over uniform sampling and separate pre-training by assigning larger weights to more informative parts of the 3D scene during sampling. Finally, the Prototype Learning scheme (fourth line) introduces interaction of LiDAR and camera encoders into pre-training and achieve the best performance.
@@ 5 Conclusion
In this paper, we propose CLAP to jointly pre-train for fusion perception in an unsupervised manner via differentiable rendering. To make joint pre-training feasible, CLAP utilizes the novel Curvature Sampling to sample more informative parts of the 3D scene for pre-training. Further, a swapping prototype prediction scheme is introduced in order to make full use of the interaction between LiDAR point clouds and camera images. To make the learnable prototypes better represent parts/semantics of objects, CLAP brings the idea of Expectation-Maximization for prototypes optimization, alongside with a gram matrix regularization term to avoid collapse. Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up. In the future, when a larger pre-training dataset is available, we expect to scale up the pre-training with CLAP for more performance improvement.
Supplementary Material
Init. F.T. mAP Vehicle Pedestrian Cyclist
0-30m 30-50m 50m- 0-30m 30-50m 50m- 0-30m 30-50m 50m-
Rand* 5% 20.48 58.03 25.22 12.98 11.62 9.75 6.97 21.55 6.83 3.11
CLAP* 22.86 +2.38 58.37 26.38 14.07 12.60 9.50 7.88 30.08 10.50 5.39
Rand 46.07 76.71 51.15 31.84 37.53 20.12 9.84 62.00 42.61 24.18
CLAP 46.88 +0.81 76.98 51.64 31.31 38.79 20.60 9.74 63.75 43.21 26.83
Rand* 100% 64.00 86.21 70.20 58.20 57.80 41.18 23.55 75.95 61.45 45.80
CLAP* 64.74 +0.74 88.14 72.59 59.13 57.37 42.24 24.22 77.11 61.91 45.63
Rand 65.03 88.18 74.23 61.75 57.32 38.90 21.96 78.07 64.32 48.16
CLAP 65.56 +0.53 87.97 72.77 62.11 58.33 40.11 21.29 78.63 64.70 47.27
Table 4: Results for transferring experiments on Once [29] dataset. CenterPoint [54] is used as the downstream detector. “Init.” indicates the initialization methodshorthands. “F.T.” indicates the number of training samples in fine-tuning stage. Overall mAP and APs for different categories within different ranges are shown in this table. “Rand*” means training the randomly initialized model with the original training iterations in OpenPCDet [42]. “Rand” indicates that we increase the number of training iterations for randomly initialized model until convergence is observed. “CLAP*” indicates that we pre-train the backbones with CLAP on NuScenes [2] and then fine-tune on Once with the original iterations in [42]. “CLAP” uses the same fine-tuning iterations as “Rand”. We use green color to highlight the performance improvement brought by CLAP. All the results are in %.
A Transferring to other datasets
We conduct further experiments to evaluate the transferring ability of CLAP. Specifically, we select LiDAR-based 3D object detector CenterPoint [54] on Once [29] for the downstream task. A 40-beam LiDAR is utilized in Once [29] to collect 15k labeled training data. We randomly sample $5\%$ and also use all of the labeled training set to train the from-scratch model until convergence is observed. Then we use pre-trained weights by CLAP on NuScenes [2] to initialize the same model and fine-tune it with the same training iterations as the randomly initialized model. Results are shown in Table 4. It can be found that pre-training by CLAP also benefits LiDAR-based 3D object detection, even in a cross-dataset setting. And if we look at the performance of “Rand*” and “CLAP*”, CLAP also accelerates the convergence in downstream task.
B Visualization of Prototype Learning
We use the model pre-trained by CLAP to infer the 3D features and assign prototypes to different LiDAR points in the 3D space. Then we use different colors to indicate different prototypes and visualize them in Figure 5. If can be found that the background road plane inside the same frame is generally assigned to the same prototype. And foreground vehicles are assigned to another prototype. This demonstrate that the proposed prototype learning scheme actually learns some level of semantics understanding in an unsupervised manner. However, as our pre-training does not incorporate any label, it can also be found that the prototype assignment has some noise, for example some of the road plane points are assigned to other prototypes.
Figure 5: Visualization of the prototype learning results. Different color indicates different prototype assignments.

### v2

@@ 4.2 Main Results
NuScenes Results. As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [58] and the best among all initialization methods. For NDS metric, UniPAD [58] only achieves comparable performance and PPKT [28] makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% . When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
Init. Level-1 Level-2 $\bar{\Delta}$
mAP mAPH mAP mAPH
Rand. 61.60 58.58 55.62 52.87 0
ALSO 62.09 59.03 56.12 53.32 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.47}}$
OCC-MAE 62.33 59.32 56.36 53.63 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.74}}$
SLidR 62.10 59.09 56.10 53.36 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.49}}$
PPKT 62.32 59.22 56.37 53.55 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.69}}$
UniPAD 61.57 58.64 55.64 52.93 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.02}}$
CLAP 62.87 59.88 56.88 54.16 $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+1.28}}$
Table 2: Results for fine-tuning on $1\%$ of training set in Waymo [45]. We first increase training iterations of from-scratch model and observe convergence (Rand.). Then we fix the training schedule to fine-tune the pre-training models. We provide an average difference $\bar{\Delta}$ to show the performance gain. All results are in %.
Waymo Results. In Table 2, we provide the average performance difference of mAP and mAPH over different difficulty levels. It can be found that CLAP achieves the best performance at convergence. Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE [34]) of previous pre-training methods. This demonstrates the effectiveness and generalization ability of CLAP.
Potential Scaling Property. As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data. The results are shown in Table 3. It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data. These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
@@
Table 3: Results on potential scaling property of CLAP . “Fine-tune Data” stands for the sample number of training set in NuScenes [3]. Results for random initialization are at convergence and fine-tuning with CLAP uses the same numbers of iterations. Results are mAP and NDS in %.
@@ 4.3 Visualizations
We use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3. The visualization results are shown in Figure 4. We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation. Furthermore, we zoom in some orange boxes. It can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
@@ 4.4 Ablation Study
We conduct ablation study to evaluate the effectiveness of different components. Results are in Table 4. The first line is seperate pre-training with UniPAD [58]. The second one is jointly pre-training using UniPAD [58] and uniform sampling guided by depth (“Memory-friendly Ray Sampling” in [58]) to address the GPU memory limitation. It can be found that using simple sampling method does not bring improvement over separate pre-training. Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by depth and separate pre-training. Finally, the Prototype Learning scheme (fourth line) learns a common feature space for segments of 3D scenes and introduces interaction of LiDAR and camera encoders, which achieves the best performance.
@@ 5 Conclusion
In this paper, we propose CLAP for unsupervised fusion joint pre-training via differentiable rendering. CLAP uses Curvature Sampling to sample more informative parts and learnable prototypes to represent parts of 3D scenes, optimized with an Expectation-Maximization approach. To further make use of the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used. Finally, a Gram Matrix Regularization loss is proposed to avoid collapse. Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up.